# MindChat: A Privacy-preserving Large Language Model for Mental Health Support

### v1

@@ Abstract
keywords:
Large language models; Mental health; Multi-agent; Privacy-preserving; Federated learning
1 Introduction
@@
Mental health is a fundamental component of overall well-being and has become a growing public health priority. Globally, approximately $4.7\%$ of the population experiences mental health disorders such as depression each year herrman2022time. The incorporation of artificial intelligence (AI) into psychotherapeutic practices presents a viable solution for broadening access to mental health assistance. Specifically, AI applications in the field of psychology cover affective computing mariani2022ai, disease diagnosis zad2021emotion, and therapeutic interventions lee2021artificial. Recent advances in large language models (LLMs), including ChatGPT achiam2023gpt, DeepSeek liu-2024-deepseek, and Qwen yang2025qwen3, has greatly broadened the range of practical AI applications. However, the responses generated by the general LLMs in the psychological counseling service scenario are often broad but lack depth and professional targeting. This limitation has prompted the emergence of a research route, that is, to improve model performance by fine-tuning on high-quality consulting datasets in specific fields. Representative efforts include SoulChat chen-etal-2023-soulchat, CPsyCounX zhang-etal-2024-cpsycoun, EmoLLM yang2024emollm, and MeChat qiu-etal-2024-smile. These efforts significantly improve the professionalism of the model in psychological counseling applications, and further emphasize the importance of high-quality emotional support datasets and effective fine-tuning strategies for the development of reliable and professional mental health LLMs.
@@
Supervised fine-tuning of LLMs for psychological counseling typically requires large-scale multi-round dialogue datasets. However, authentic counseling conversations are scarce and often inaccessible due to privacy considerations, which greatly increases the difficulty of data collection Liu_ACL_20. To meet this challenge, prior work explores generating training data by prompting LLMs with seed instructions and partial dialogue fragments, leveraging their capacity to complete and expand consultative dialogue zheng-etal-2023-augesc. Techniques like cognitive restructuring xiao-etal-2024-healme and reasoning-augmented prompting zhang-etal-2024-escot are introduced to improve the authenticity of generated dialogues and their alignment with therapeutic principles. Nevertheless, most current approaches adopt a one-pass generation workflow, where data quality is ensured through subsequent filtering or manual revision, rather than iterative optimization through dynamic feedback mechanisms. The fine-tuning strategies of LLMs can be generally categorized into full fine-tuning and parameter-efficient fine-tuning (PEFT). Full fine-tuning requires updating all model parameters during the training process, which puts forward considerable requirements in terms of data volume and computing resources. In contrast, PEFT only adjusts a small subset of parameters, which can still achieve competitive performance while significantly reducing the cost of training ding2023parameter. Representative PEFT methods include Low-Rank Adaptation (LoRA) hu2022lora, P-Tuning liu-etal-2022-p, and Adapter Tuning chen2023hadamard, among which LoRA gains widespread adoption for its favorable trade-off between computational efficiency and task-specific performance.
@@
Dialogues in psychological counseling often involve highly sensitive personal information, making their use for model training a source of significant ethical and security concerns. Additionally, such data are typically distributed across hospitals, counseling clinics, or online platforms, showing decentralization characteristics, while local institutions need to ensure that private records are not leaked externally. These constraints hinder centralized data aggregation. Addressing these challenges requires a training paradigm that supports decentralized collaboration with low communication overhead, alongside robust privacy protection mechanisms. Privacy risks related to LLMs have attracted growing attention in recent studies. For instance, the combination of jailbreaking attacks and chain-of-thought reasoning can induce models like ChatGPT to disclose confidential personal information li-etal-2023-multi-step. From a data security standpoint, incorporating differential privacy (DP) noise into the training process serves as an effective means of mitigating privacy breaches YAO2024100211. While existing studies explore privacy protection in general LLMs wu-etal-2023-depn and federated learning (FL) is applied to train domain-specific LLMs on distributed data yue2023fedjudge, a psychological LLM that comprehensively addresses the privacy protection of training data remains under-explored.
@@
2 Related works
2.1 Data construction for LLMs
The scarcity of high-quality data poses a major challenge to the advancement of LLMs. In order to alleviate this issue, data synthesis and augmentation emerge as effective strategies for enriching training corpus wang2024survey. For instance, ChatGPT is employed to generate multi-round dialogues to improve the performance of models such as LLaMA xu-etal-2023-baize. In the medical domain, HuatuoGPT zhang-etal-2023-huatuogpt demonstrates that ChatGPT-distilled dialogues, when combined with real-world clinical data, can effectively support supervised fine-tuning of medical consultation LLMs.
Similarly, in the field of mental health applications, researchers enrich datasets by synthesizing multi-turn counseling dialogues from limited sources. Early efforts focus on expanding single-turn data: SMILECHAT qiu-etal-2024-smile uses ChatGPT to rewrite queries into empathetic multi-turn communication via prompt-based extensions. On this basis, SoulChatCorpus chen-etal-2023-soulchat strengthens empathy constraints in prompting and applies manual proofreading to produce a large-scale Chinese dataset, emphasizing supportive behaviors such as active listening and emotional validation. Despite their scale and empathetic design, these approaches often lack grounding in professional counseling knowledge, which limits their therapeutic authenticity. Subsequent work seeks to ground synthetic dialogues in domain expertise. CPsyCounD zhang-etal-2024-cpsycoun constructs dialogues directly from psychological counseling reports using a two-phase Memo2Demo pipeline, explicitly embedding counseling principles into synthetic interactions. PsyDTCorpus xie2025psydt further explores this direction by leveraging dynamic one-shot learning with GPT-4 to capture counselor linguistic styles and therapy techniques, generating personalized multi-turn dialogues conditioned on client personality. More recently, domain-specific scenarios are explored. For instance, PeConv zhao2025parental targets parent–child emotional support by integrating child emotion coaching theory into a human–machine collaborative framework, and represents an early effort toward constructing a Chinese dialogue dataset for parental counseling assistance.
@@ 2.1 Data construction for LLMs
2.2 Privacy-preserving LLMs for mental health
@@
Recent studies explore adapting LLMs to mental health support scenarios through fine-tuning on domain-specific multi-turn dialogue data, as exemplified by SoulChat chen-etal-2023-soulchat, CPsyCounX zhang-etal-2024-cpsycoun, EmoLLM yang2024emollm, and MeChat qiu-etal-2024-smile. These works demonstrate that incorporating counseling-oriented interaction histories enhances empathy and response quality in psychological LLMs. Despite their effectiveness, these methods all use the centralized training mode. Such a prerequisite inevitably raises privacy concerns, as mental health data is highly sensitive and subject to strict regulations such as HIPAA and GDPR, which greatly restrict the applicability of these fine-tuning approaches in settings where data sharing across institutions or individuals is not permitted. Moreover, LLMs exhibit a tendency to memorize training data, resulting further risks of sensitive information being reconstructed through adversarial or model inversion attacks Feldman-2020-memory.
@@
In order to mitigate these risks, existing efforts on privacy-preserving LLMs mainly focus on data preprocessing and model-level defenses. At the data level, anonymization techniques, such as removing or masking personal identification information, are commonly employed to reduce privacy exposure zuo-2021-ano1, zhou-2022-ano2. Unlike anonymization, which may alter the original content of the data, model-level approaches preserve data integrity while providing privacy protection during training. Among these, FL serves as a promising paradigm for scenarios where sensitive data is distributed across multiple parties, enabling collaborative model training without centralizing raw data chen-2024-fl.
@@
In mental health domain, FL has been explored for problems including mental health sentiment forecasting ahsan2024privacy and depression detection from multilingual textual data khalil2024federated. These studies demonstrate the worth of FL in reducing data exposure risks, yet they primarily focus on classification tasks rather than generative language modeling. More recent work extends FL to LLMs for mental health applications. FedMentalCare sarwar2025fedmentalcare presents a federated framework with PEFT for mental health analysis, demonstrating scalability and reduced computational overhead. Nevertheless, this line of research does not incorporate formal privacy protection mechanisms, nor does it consider multi-turn counseling dialogues that are essential for modeling therapeutic interactions. To further defend against privacy leakage through model parameters, DP is often integrated into federated systems through central or local mechanisms wang-2024-central, youn-2023-local. Building on these insights, the present work introduces a privacy-preserving framework that combines FL with local DP to fine-tune psychological LLMs on multi-turn counseling dialogues, reducing the risk of data reconstruction while maintaining training efficiency in distributed mental health scenarios.
3 Methodology
@@
3.1 Mental health support data construction
3.1.1 Seed data
The situation of individuals seeking psychological support play a crucial role in generating realistic mental health dialogues. Motivated by zhang-etal-2024-escot, the seed dataset is constructed using approximately 1111k situation texts related to psychological seekers, collected from the publicly accessible online counseling platforms Yixinli and Jiandanxinli. The corpus covers a broad spectrum of psychological and behavioral areas, including emotional and relationship management, self-awareness and individual growth, stress and anxiety relief, mental health maintenance, and workplace adjustment. Collectively, these diverse and thematically comprehensive data provide a solid foundation for simulating real psychological counseling scenarios. However, the raw data exhibit gaps in key contextual information and often contain overly colloquial expressions. For this reason, we use the GLM-4-Plus model to clean and enrich the original materials. The resultant individual situation representation comprises three core elements: Character, Plight, and Demand, as illustrated in Fig. 1.
Figure 1: The personal situation of seekers.
@@ 3.1.1 Seed data
3.1.2 Dialogue construction process
We propose a multi-agent architecture comprising six specialized agents: Extractor, Seeker, Supporter, Evaluator, Corrector, and Manager, integrated with a dual-loop feedback mechanism to synthesize emotional support dialogues, as shown in Fig. 2. Detailed introductions of these agents are provided in A.1, while the corresponding prompt templates used to instantiate each agent are included in A.2.
3.2 The training framework of MindChat
MindChat is trained within a parameter-efficient and privacy-preserving FL framework, as illustrated in Fig. 3. The design integrates LoRA for efficient on-device fine-tuning and client-level DP to safeguard sensitive mental health dialogues during collaborative training.
3.2.1 Efficient fine-tuning
@@
where $\Delta\bm{W}_{i}^{t}=\bm{B}_{i}^{t}\bm{A}_{i}^{t}$ denotes the LoRA-induced update. Since $r\ll\min(d,k)$, the scale of trainable parameters is significantly reduced, resulting in lower computation and communication costs.
3.2.2 Privacy protection federated training
To protect sensitive counseling data during federated training, client-level $(\epsilon,\delta)$-DP is enforced on the model updates transmitted to the server, where $\epsilon$ controls the privacy budget and $\delta$ denotes the probability of privacy failure. Each client $C_{i}$ holds a frozen base model $\bm{W}^{0}$ and, at communication round $t$, generates its update by optimizing only the low-rank adapter matrices $\bm{A}_{i}^{t}$ and $\bm{B}_{i}^{t}$. Let $\bm{\theta}_{i}^{t}=\{\bm{A}_{i}^{t},\bm{B}_{i}^{t}\}$ denote the collection of all trainable parameters in the LoRA adapter matrices on client $C_{i}$ at round $t$, and let $\bm{\theta}_{s}^{t-1}$ denote the corresponding global LoRA parameters received from the server. After local training, the client computes its LoRA parameter update
$\Delta\bm{\theta}_{i}^{t}=\bm{\theta}_{i}^{t}-\bm{\theta}_{s}^{t-1}.$ (2)
To control the magnitude of the client update, we apply $L_{2}$-norm clipping to $\Delta\bm{\theta}_{i}^{t}$ with a pre-specified threshold $C$:
@@ 3.2.2 Privacy protection federated training
Upon receiving the privacy-preserving LoRA parameters from participating clients, the central server aggregates them using the FedAvg algorithm mcmahan2017communication. Specifically, at round $t$, the global LoRA parameters are updated as
@@
where $|D_{i}|$ denotes the amount of local training samples on client $C_{i}$. The aggregated global LoRA parameters $\bm{\theta}_{s}^{t}$ are then broadcast to all clients to initialize the next round of training. After the final communication round, the resulting LoRA parameters $\bm{\theta}_{s}^{t}=\{\bm{A}_{s}^{t},\bm{B}_{s}^{t}\}$ are combined with the frozen base model to form the complete model weight matrix
$\bm{W}_{s}^{t}=\bm{W}^{0}+\bm{B}_{s}^{t}\bm{A}_{s}^{t}.$ (7)
Table 1: Details of the five dimensions in data evaluation.
@@ 3.3 Evaluation methods
The quality of conversational corpora directly influences the performance of models trained on it. In research on evaluating synthetic data for mental health assistance, current evaluation dimensions are typically supporter-centered, focusing on the reasonable response of the supporter zheng-etal-2024-self. However, in reality, the effectiveness of psychological consultations is ultimately determined by the feelings of the seeker. Inspired by xiao-etal-2024-healme, Hill_helping_20, this paper proposes five evaluation dimensions, Professionalism (Pro.), Helpfulness (Hel.), Guidance (Gui.), Emotion (Emo.), and Trust (Tru.), to comprehensively assess the generated data from both the perspectives of the seeker and supporter. These dimensions are thoroughly explained in Table 1.
Particularly, CpsyCounE zhang-etal-2024-cpsycoun, a dataset designed to evaluate the psychological counseling capabilities of LLMs, is utilized to assess MindChat and baselines. According to zhang-etal-2024-cpsycoun, the evaluation is performed from four dimensions: Comprehensiveness, Professionalism, Authenticity, and Safety. Comprehensiveness refers to the extent to which the dialogue covers the background of users and psychological concerns, Professionalism indicates the proficiency of the model in psychological counseling, Authenticity reflects the degree to which the dialogue aligns with real-world scenarios and Safety denotes the level of privacy protection provided to the user. In addition, we adopt ROUGE-1 and cosine similarity as proxies for data memorization, and further assess privacy guarantees through membership inference attacks(MIAs), quantified using both ROC AUC murakonda2021quantifying and PR AUC shokri2017membership. These metrics collectively reflect the strength of privacy protection under DP.
4 Experiments

### v2

# Towards Privacy-Preserving Mental Health Support with Large Language Models

@@ Abstract
I Introduction
@@
Mental health is a fundamental component of overall well-being and has become a growing public health priority. Globally, approximately $4.7\%$ of the population experiences mental health disorders such as depression each year [16]. The incorporation of artificial intelligence (AI) into psychotherapeutic practices presents a viable solution for broadening access to mental health assistance. Specifically, AI applications in the field of psychology cover affective computing [15], disease diagnosis [38], and therapeutic interventions [22]. Recent advances in large language models (LLMs), including ChatGPT [1], DeepSeek [26], and Qwen [42], has greatly broadened the range of practical AI applications. However, the responses generated by the general LLMs in the psychological counseling service scenario are often broad but lack depth and professional targeting. This limitation has prompted the emergence of a research route, that is, to improve model performance by fine-tuning on high-quality consulting datasets in specific fields. Representative efforts include SoulChat [8], CPsyCounX [47], EmoLLM [43], and MeChat [32]. These efforts significantly improve the professionalism of the model in psychological counseling applications, and further emphasize the importance of high-quality emotional support datasets and effective fine-tuning strategies for the development of reliable and professional mental health LLMs.
@@
Supervised fine-tuning of LLMs for psychological counseling typically requires large-scale multi-round dialogue datasets. However, authentic counseling conversations are scarce and often inaccessible due to privacy considerations, which greatly increases the difficulty of data collection [27]. To meet this challenge, prior work explores generating training data by prompting LLMs with seed instructions and partial dialogue fragments, leveraging their capacity to complete and expand consultative dialogue [51]. Techniques like cognitive restructuring [39] and reasoning-augmented prompting [49] are introduced to improve the authenticity of generated dialogues and their alignment with therapeutic principles. Nevertheless, most current approaches adopt a one-pass generation workflow, where data quality is ensured through subsequent filtering or manual revision, rather than iterative optimization through dynamic feedback mechanisms. Based on the constructed data, the training of LLMs typically relies on fine-tuning strategies, which can be generally categorized into full fine-tuning and parameter-efficient fine-tuning (PEFT). Full fine-tuning requires updating all model parameters during the training process, which puts forward considerable requirements in terms of data volume and computing resources. In contrast, PEFT only adjusts a small subset of parameters, which can still achieve competitive performance while significantly reducing the cost of training [11]. Representative PEFT methods include Low-Rank Adaptation (LoRA) [18], P-Tuning [28], and Adapter Tuning [9], among which LoRA gains widespread adoption for its favorable trade-off between computational efficiency and task-specific performance.
@@
Dialogues in psychological counseling often involve highly sensitive personal information, making their use for model training a source of significant ethical and security concerns. Additionally, such data are typically distributed across hospitals, counseling clinics, or online platforms, showing decentralization characteristics, while local institutions need to ensure that private records are not leaked externally. These constraints hinder centralized data aggregation. Addressing these challenges requires a training paradigm that supports decentralized collaboration with low communication overhead, alongside robust privacy protection mechanisms. Meanwhile, privacy risks related to LLMs have attracted growing attention in recent studies. For instance, the combination of jailbreaking attacks and chain-of-thought reasoning can induce models like ChatGPT to disclose confidential personal information [24]. From a data security standpoint, incorporating differential privacy (DP) noise into the training process serves as an effective means of mitigating privacy breaches [44]. While existing studies explore privacy protection in general LLMs [37] and federated learning (FL) is applied to train domain-specific LLMs on distributed data [10], a psychological LLM that comprehensively addresses the privacy protection of training data remains under-explored.
@@
II Related Works
II-A Data Construction for LLMs
The scarcity of high-quality data poses a major challenge to the advancement of LLMs. In order to alleviate this issue, data synthesis and augmentation emerge as effective strategies for enriching training corpus [5]. For instance, ChatGPT is employed to generate multi-round dialogues to improve the performance of models such as LLaMA [41]. In the medical domain, HuatuoGPT [48] demonstrates that ChatGPT-distilled dialogues, when combined with real-world clinical data, can effectively support supervised fine-tuning of medical consultation LLMs.
Similarly, in the field of mental health applications, researchers enrich datasets by synthesizing multi-turn counseling dialogues from limited sources. Early efforts focus on expanding single-turn data: SMILECHAT [32] uses ChatGPT to rewrite queries into empathetic multi-turn communication via prompt-based extensions. On this basis, SoulChatCorpus [8] strengthens empathy constraints in prompting and applies manual proofreading to produce a large-scale Chinese dataset, emphasizing supportive behaviors such as active listening and emotional validation. Despite their scale and empathetic design, these approaches often lack grounding in professional counseling knowledge, which limits their therapeutic authenticity. Subsequent work seeks to ground synthetic dialogues in domain expertise. CPsyCounD [47] constructs dialogues directly from psychological counseling reports using a two-phase Memo2Demo pipeline, explicitly embedding counseling principles into synthetic interactions. PsyDTCorpus [40] further explores this direction by leveraging dynamic one-shot learning with GPT-4 to capture counselor linguistic styles and therapy techniques, generating personalized multi-turn dialogues conditioned on client personality. More recently, domain-specific scenarios are explored. For instance, PeConv [50] targets parent–child emotional support by integrating child emotion coaching theory into a human–machine collaborative framework, and represents an early effort toward constructing a Chinese dialogue dataset for parental counseling assistance.
@@ II-A Data Construction for LLMs
II-B Privacy-preserving LLMs for Mental Health
@@
Recent studies explore adapting LLMs to mental health support scenarios through fine-tuning on domain-specific multi-turn dialogue data, as exemplified by SoulChat [8], CPsyCounX [47], EmoLLM [43], and MeChat [32]. These works demonstrate that incorporating counseling-oriented interaction histories enhances empathy and response quality in psychological LLMs. Despite their effectiveness, these methods all use the centralized training mode. Such a prerequisite inevitably raises privacy concerns, as mental health data is highly sensitive and subject to strict regulations such as HIPAA and GDPR, which greatly restrict the applicability of these fine-tuning approaches in settings where data sharing across institutions or individuals is not permitted. Moreover, LLMs exhibit a tendency to memorize training data, resulting further risks of sensitive information being reconstructed through adversarial or model inversion attacks [14].
@@
In order to mitigate these risks, existing efforts on privacy-preserving LLMs mainly focus on data preprocessing and model-level defenses. At the data level, anonymization techniques, such as removing or masking personal identification information, are commonly employed to reduce privacy exposure [54]. Unlike anonymization, which may alter the original content of the data, model-level approaches preserve data integrity while providing privacy protection during training. Among these, FL serves as a promising paradigm for scenarios where sensitive data is distributed across multiple parties, enabling collaborative model training without centralizing raw data [7].
@@
In mental health domain, FL has been explored for situations including facial emotion recognition [6] and depression detection from multilingual textual data [20]. These studies demonstrate the worth of FL in reducing data exposure risks, yet they primarily focus on classification tasks rather than generative language modeling. More recent work extends FL to LLMs for mental health applications. FedMentalCare [33] presents a federated framework with PEFT for mental health analysis, demonstrating scalability and reduced computational overhead. Nevertheless, this line of research does not incorporate formal privacy protection mechanisms, nor does it consider multi-turn counseling dialogues that are essential for modeling empathetic interactions. To further defend against privacy leakage through model parameters, DP is often integrated into federated systems through central or local mechanisms [19]. Building on these insights, the present work introduces a privacy-preserving framework that combines FL with local DP to fine-tune psychological LLMs on multi-turn counseling dialogues, reducing the risk of data reconstruction while maintaining training efficiency in distributed mental health scenarios.
III Methodology
Figure 1: The personal situation of seekers.
@@ III Methodology
III-A Mental Health Support Data Construction
III-A1 Seed data
The situation of individuals seeking psychological support play a crucial role in generating realistic mental health dialogues. Motivated by [49], the seed dataset is constructed using approximately 1111k situation texts related to psychological seekers, collected from the publicly accessible online counseling platforms Yixinli and Jiandanxinli. The corpus covers a broad spectrum of psychological and behavioral areas, including emotional and relationship management, self-awareness and individual growth, stress and anxiety relief, mental health maintenance, and workplace adjustment. Collectively, these diverse and thematically comprehensive data provide a solid foundation for simulating real psychological counseling scenarios. However, the raw data exhibit gaps in key contextual information and often contain overly colloquial expressions. For this reason, we use the GLM-4-Plus model to clean and enrich the original materials. The resultant individual situation representation comprises three core elements: Character, Plight, and Demand, as illustrated in Fig. 1.
@@ III-A1 Seed data
III-A2 Dialogue construction process
We propose a multi-agent architecture comprising six specialized agents: Extractor, Seeker, Supporter, Evaluator, Corrector, and Manager, integrated with a dual-loop feedback mechanism to synthesize emotional support dialogues, as shown in Fig. 2. Detailed introductions of these agents are provided in Appendix A, while the corresponding prompt templates used to instantiate each agent are included in Appendix B.
III-B The Training Framework
@@
The model of MindChat is trained within a parameter-efficient and privacy-preserving FL framework, as illustrated in Fig. 3. The design integrates LoRA for efficient on-device fine-tuning and client-level DP to safeguard sensitive mental health dialogues during collaborative training.
III-B1 Efficient fine-tuning
@@
where $\Delta\bm{W}_{i}^{t}=\bm{B}_{i}^{t}\bm{A}_{i}^{t}$ denotes the LoRA-induced update. Since one can obtain that $r\ll\min(d,k)$, the scale of trainable parameters is significantly reduced, resulting in lower computation and communication costs.
III-B2 Privacy protection federated training
To protect sensitive counseling data during federated training, client-level $(\epsilon,\delta)$-DP is enforced on the model updates transmitted to the server, where $\epsilon$ controls the privacy budget and $\delta$ denotes the probability of privacy failure. Each client $C_{i}$ holds a frozen base model $\bm{W}^{0}$ and, at communication round $t$, generates its update by optimizing only the low-rank adapter matrices $\bm{A}_{i}^{t}$ and $\bm{B}_{i}^{t}$. Let $\bm{\theta}_{i}^{t}=\{\bm{A}_{i}^{t},\bm{B}_{i}^{t}\}$ denote the collection of all trainable parameters in the LoRA adapter matrices on client $C_{i}$ at round $t$, and let $\bm{\theta}_{s}^{t-1}$ denote the corresponding global LoRA parameters received from the server. After local training, the client computes its LoRA parameter update as $\Delta\bm{\theta}_{i}^{t}=\bm{\theta}_{i}^{t}-\bm{\theta}_{s}^{t-1}$. To control the magnitude of the client update, we apply $L_{2}$-norm clipping to $\Delta\bm{\theta}_{i}^{t}$ with a pre-specified threshold $C$:
@@ III-B2 Privacy protection federated training
Upon receiving the privacy-preserving LoRA parameters from participating clients, the central server aggregates them using the FedAvg algorithm [30]. Specifically, at round $t$, the global LoRA parameters are updated as
@@
where $|D_{i}|$ denotes the amount of local training samples on client $C_{i}$. The aggregated global LoRA parameters $\bm{\theta}_{s}^{t}$ are then broadcast to all clients to initialize the next round of training. After the final communication round, the resulting LoRA parameters $\bm{\theta}_{s}^{t}=\{\bm{A}_{s}^{t},\bm{B}_{s}^{t}\}$ are combined with the frozen base model to form the complete model weight matrix $\bm{W}_{s}^{t}=\bm{W}^{0}+\bm{B}_{s}^{t}\bm{A}_{s}^{t}$.
TABLE I: Details of the Five Dimensions in Data Evaluation.
@@ III-C Evaluation Methods
The quality of conversational corpora directly influences the performance of models trained on it. In research on evaluating synthetic data for mental health assistance, current evaluation dimensions are typically supporter-centered, focusing on reasonable response of the supporter [52]. However, in reality, effectiveness of psychological consultations is ultimately determined by feelings of the seeker. Inspired by [39, 17], this paper proposes five evaluation dimensions, Professionalism (Pro.), Helpfulness (Hel.), Guidance (Gui.), Emotion (Emo.), and Trust (Tru.), to comprehensively assess the generated data from both seeker and supporter perspectives. These dimensions are thoroughly explained in Table I.
Particularly, CpsyCounE [47], a dataset designed to evaluate the psychological counseling capabilities of LLMs, is utilized to assess MindChat and baselines. According to [47], the evaluation is performed from four dimensions: Comprehensiveness, Professionalism, Authenticity, and Safety. Comprehensiveness refers to the extent to which the dialogue covers the background of users and psychological concerns, Professionalism indicates the proficiency of the model in psychological counseling, Authenticity reflects the degree to which the dialogue aligns with real-world scenarios and Safety denotes the level of privacy protection provided to the user. In addition, we adopt ROUGE-1 and cosine similarity as proxies for data memorization, and further assess privacy guarantees through membership inference attacks (MIAs), quantified using both ROC AUC [31] and PR AUC [35]. These metrics collectively reflect the strength of privacy protection under DP.
IV Experiments